Assuming no change in temperature and pressure, Calculate the volume of O2 (in liters) required for the complete combustion of 14.9 L of butane (C4H10):

                                        C4H10(g) + O2(g)       →  CO2(g) + H2O(l )    

1
Expert's answer
2021-12-15T10:17:21-0500

Q279316


Assuming no change in temperature and pressure, Calculate the volume of O2 (in liters) required for the complete combustion of 14.9 L of butane (C4H10):

                                        C4H10(g) + O2(g)       →  CO2(g) + H2O(l )   



Solution :


 

"C_4H_{10} (g) + O_2(g) \u2192 CO_2(g) + H_2O(l )"

Let us first balance this reaction.


there is 4 C on left side and 1 C on right side. If we write 4 in front of CO 2, then both side will contain same number of C.


"C_4H_{10} (g) + O_2(g) → 4 \ CO_2(g) + H_2O(l )"


There is 10 H on left side and 2 H on right side. For balancing H we will write 5 in front of H2O.


"C_4H_{10} (g) + O_2(g) → 4 \ CO_2(g) + 5 \ H_2O(l )"


Lastly we will balance the O on both the side.

There is 2 O on left side and 13 O on right side.

A proper fractional number must be written in front of O2, so that both side contains same number of O.

If we write 13/2 in front of O2, then both side will contain same number of O which is 13.


 

"C_4H_{10} (g) +\frac{13}{2} O_2(g) → 4 \ CO_2(g) + 5 \ H_2O(l )"


In the balanced chemical reaction the coefficient can't be fraction.

multiply whole reaction by 2 for removing the fraction.



"2 * [ \ C_4H_{10} (g) +\frac{13}{2} O_2(g) → 4 \ CO_2(g) + 5 \ H_2O(l ) \ ]"


"2 \ C_4H_{10} (g) \ + \ 13 \ O_2(g) → 8 \ CO_2(g) + 10 \ H_2O(l )"


Now the reaction is balanced for all the elements




"\textcolor{darkblue}{2 \ C_4H_{10} (g) \ + \ 13 \ O_2(g) → 8 \ CO_2(g) + 10 \ H_2O(l ) }"

Now we will find the volume of H2O in Liters formed from combustion of 14.9 L of butane ( C4H10)


In the given reaction you can see that 2 volume of C4H10 reacts with 13 volume of O2 to give 8 volume of CO2 and 10 volume of H2O.


So volume of H2O formed on combustion of 14.9 L of butane is



"=14.9 \ \cancel{L \ C_4H_{10}} * \frac{10 \ L \ H_2O }{2 \ \cancel{L \ C_4H_{10}} }"


"= 74.5 \ L \ H_2O"


Hence 74.5 L of H2O will be formed on complete combustion of 14.9 L of butane.